A 75 kg patient needs a medication X in the amount of 0.2 mg/kg/day. The medication comes in a liquid solution of 50 mg for every 1000 ml. How many ml of this solution will this patient need in a day? If this solution is being infused at 60 ml/hr, how long with the transfusion take?
Answer #1

ANSWER

* ml of solution patient needed in a day.

       Formula

           Amount of solution needed = Amount desired (D) / Amount in hand(H) * Quantity(Q)

              Weight of the patient = 75 kg.

               Amount desired (D) = 0.2mg/kg/day.

             So amount required for this patient = 0.2 * 75

= 15 mg / day.

   Amount in hand (H) = 50mg

   Quantity = 1000 ml.      

   Amount (ml) of solution needed in a day = D / H *Q

   = 15 / 50 *1000

       = 0.3 * 1000

      = 300 ml.

* Time taken for transfusion.

    * Solution is infused in 60 ml /hr.

Amount of solution need to be infused = 300 ml.

      Time taken for transfusion = Amount of solution need to be infused / solution infused per hour.      

   = 300 / 60

= 5 hr.